### v1

@@ E.1 The Algorithm
    The probability of event (a) is $\frac{1}{p^{t}}$. Now, conditioned on event (a). Since the $v_{1},\ldots,v_{t}$ and $u$ are all independent random variables, the conditional probability of event (b) is $\frac{1}{p^{t}}$. We thus have that $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{t}}\cdot\frac{1}{p^{t}}=\frac{1}{p^{2t}}$.
    Now, $\mathbb{E}[I_{\chi\to b+u}]=\Pr[\chi\in C_{u}(b)]=\frac{1}{p^{t}}$ by part (1). Similarly, $\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{t}}$. We thus have that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=\frac{1}{p^{2t}}-\frac{1}{p^{t}}\cdot\frac{1}{p^{t}}=0$.
    Case 2: $\chi,\chi^{\prime}$ are both nonzero, distinct and linearly independent.
@@
    $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$ and $\langle\chi^{\prime},v_{i}\rangle=b_{i}+u_{i}$. Fix $i\in[t]$ and $u_{i}$. Since $\chi$ and $\chi^{\prime}$ are linearly independent, the simultaneous solutions of the linear equations $\langle\chi,v_{i}\rangle=b_{i}+u_{i}$ and $\langle\chi^{\prime},v_{i}\rangle=b_{i}+u_{i}$ span an affine subspace of $\mathbb{Z}_{p}^{n}$ of co-dimension $2$. Thus, for each $i$, the probability that $v_{i}$ belongs to this affine subspace is $\frac{1}{p^{2}}$. Since $u_{1},\ldots,u_{t},v_{1},\ldots,v_{t}$ are all independent, we have that $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{2t}}$.
    Now, for the same reason as in the previos case, $\mathbb{E}[I_{\chi\to b+u}]=\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{t}}$. we thus have that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=\frac{1}{p^{2t}}-\frac{1}{p^{t}}\cdot\frac{1}{p^{t}}=0$.
@@
Finally we are ready to prove Lemma E.5.
Proof E.17 (Proof of Lemma E.5).
@@
Since $f$ is $\epsilon$-far from any $s$-sparse Boolean function with $\epsilon\leq\frac{1}{8(p^{2}s)^{p-1}}$, so by Theorem 1.4, it is $\frac{\epsilon}{\sqrt{2}}$-far from any $s$-sparse function. We will prove the lemma in two cases.
@@
In this case from Theorem E.13 (part (2)) implies that with probability at least $\frac{19}{20}$, at least $s+1$ buckets will contain elements from $A_{1}$. Therefore, the weights of those buckets will be at least $\tau$, and hence it follows.
Case 2: $|A_{1}|\leq s$ $\mathsf{wt}(A_{2})\geq\epsilon^{2}/2$.
Let us define the random variables $Y_{b}=\mathsf{wt}(C_{u}(b)\cap A_{2})$ for each $b$. Then $Y_{b}=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$, where $I_{\chi\to b+u}$ is the indicator random variable which takes the value $1$ if $\chi\in C_{u}(b)$ and 00 otherwise.
@@
$\mathbb{E}[Y_{b}]=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\Pr[\chi\in C_{u}(b)]=\frac{1}{p^{t}}\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\geq\frac{\epsilon^{2}}{2p^{t}},$ (E.2)
where the second equality follows from the fact that $\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\geq(\frac{\epsilon}{\sqrt{2}})^{2}$ and from Theorem E.13 (part (1)) where we show that $\Pr[\chi\in C_{u}(b)]=\frac{1}{p^{t}}$.
Before we apply Chebyshev’s inequality to upper bound the probability that a bucket has small weight we upper bound the variance of $Y_{b}$. We can write $Y_{b}=\sum_{\chi\in A_{2}}Y_{\chi,b}$, where $Y_{\chi,b}=|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$. Note that $0\leq Y_{b}\leq\tau$, $\mathbb{E}[Y_{b}]>0$ (by Equation E.2) and $\mathsf{cov}(Y_{\chi,b},Y_{\chi^{\prime},b})=0$ (by Theorem E.13, where $\chi,\chi^{\prime}\in A_{2}$). So, by Lemma E.15,
@@
$\Pr[Y_{b}\leq(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]]$ $\leq\frac{\operatorname{Var}[Y_{b}]}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]^{2}}\leq\frac{\tau\mathbb{E}[Y_{b}]}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]^{2}}=\frac{\tau}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]}\leq\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\cdot\epsilon^{2}}.$
Since the total number of buckets is $p^{t}$, therefore the expected number of buckets with weight $\leq(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $p^{t}\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}$. Thus by Markov’s inequality, with probability at least $\frac{9}{10}$, the number of buckets with weight more that $(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}$.
@@
$(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]\geq\frac{1}{20}\cdot\frac{\epsilon^{2}}{2p^{t}}=\frac{\epsilon^{2}}{40p^{t}}\geq\tau.$
@@
It thus follows that with probability at least $\frac{9}{10}$, the number of buckets with weight more that $(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}$.
@@
We now finish by showing that $10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}\leq p^{t}-(s+1)$.
$10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}\leq p^{t}-(s+1)$
@@
$\Longleftarrow\tau\leq\frac{\epsilon^{2}}{p^{t}}\cdot\biggl{(}1-\frac{s+1}{p^{t}}\biggr{)}\cdot\left(\frac{19}{20}\right)^{2}\cdot\frac{1}{20}$
$\Longleftarrow\tau\leq\frac{\epsilon^{2}}{p^{t}}\cdot\frac{1}{2}\cdot\frac{1}{20}=\frac{\epsilon^{2}}{40p^{t}}.$

### v2

@@ 6.1 The Algorithm
    The probability of event (a) is $\frac{1}{\Delta^{t}}$. Now, conditioned on event (a). Since the $v_{1},\ldots,v_{t}$ and $u$ are all independent random variables, the conditional probability of event (b) is $\frac{1}{\Delta^{t}}$. We thus have that $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{t}}\cdot\frac{1}{\Delta^{t}}=\frac{1}{\Delta^{2t}}$.
    Now, $\mathbb{E}[I_{\chi\to b+u}]=\Pr[\chi\in C_{u}(b)]=\frac{1}{\Delta^{t}}$ by part (1). Similarly, $\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{t}}$. We thus have that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=\frac{1}{\Delta^{2t}}-\frac{1}{\Delta^{t}}\cdot\frac{1}{\Delta^{t}}=0$.
    Subcase 2: ($\lambda-1$) is not invertible in $\mathbb{Z}_{\Delta}$.
    Then $\lambda-1$ is a zero divisor in $\mathbb{Z}_{\Delta}$. It follows that $b_{i}+u_{i}=m_{i}$, for some $m_{i}$ such that $m_{i}$ divides $\Delta$, for all $i\in[t]$. Observe that $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if
    1. (a)
      $b_{i}+u_{i}=m_{i}$ for all $i\in[t]$, i.e., $u_{i}=m_{i}-b_{i}$ for all $i\in[t]$, and
    2. (b)
      $\forall i\in[t],\langle\gamma_{1},\beta_{i}\rangle=b_{i}+u_{i}$.
    Clearly, $\Pr[u_{i}=m_{i}-b_{i}\ \forall i\in[t]]=\frac{1}{\Delta^{t}}$, since the value of $\lambda$ determines that of $m_{i}$’s. Since the $\beta_{1},\ldots,\beta_{t}$ and $u$ are all independent random variables, $\Pr[\forall i\in[t],\langle\chi,\beta_{i}\rangle=b_{i}+u_{i}|u_{i}=m_{i}-b_{i}]=\frac{1}{\Delta^{t}}$. Therefore, $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{t}}\cdot\frac{1}{\Delta^{t}}=\frac{1}{\Delta^{2t}}$.
    Now, $\mathbb{E}[I_{\chi\to b+u}]=\Pr[\chi\in C_{u}(b)]=\frac{1}{\Delta^{t}}$. Similarly, it follows that $\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{t}}$. We thus have $\operatorname{Cov}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{2t}}-\frac{1}{\Delta^{2t}}=0$.
    Case 2: $\chi,\chi^{\prime}$ are both nonzero, distinct and one is not a multiple of the other.
@@
    $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$ and $\langle\chi^{\prime},v_{i}\rangle=b_{i}+u_{i}$. Fix $i\in[t]$ and $u_{i}$. Since $\chi^{\prime}$ is not a multiple of $\chi$, the bucket $\chi^{\prime}$ belongs to is not determined by that of $\chi$. Thus, for each $i$, the probability that $v_{i}$ belongs to this bucket is $\frac{1}{\Delta^{2}}$. Since $u_{1},\ldots,u_{t},v_{1},\ldots,v_{t}$ are all independent, we have that $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{2t}}$.
    Now, for the same reason as in the previous case, $\mathbb{E}[I_{\chi\to b+u}]=\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{\Delta^{t}}$. we thus have that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=\frac{1}{\Delta^{2t}}-\frac{1}{\Delta^{t}}\cdot\frac{1}{\Delta^{t}}=0$.
@@
Finally we are ready to prove Lemma 6.5.
Proof of Lemma 6.5.
@@
Since $f$ is $\epsilon$-far from any $s$-sparse Boolean function with $\epsilon\leq\frac{1}{8(2s)^{\varphi(\Delta)}}$, so by Theorem 1.4, it is $\frac{\epsilon}{\sqrt{2}}$-far from any $s$-sparse function. We will prove the lemma in two cases.
@@
In this case from Theorem 6.10 (part (2)) implies that with probability at least $\frac{19}{20}$, at least $s+1$ buckets will contain elements from $A_{1}$. Therefore, the weights of those buckets will be at least $\tau$, and hence it follows.
Case 2: $|A_{1}|\leq s$ & $\mathsf{wt}(A_{2})\geq\epsilon^{2}/2$.
Let us define the random variables $Y_{b}=\mathsf{wt}(C_{u}(b)\cap A_{2})$ for each $b$. Then $Y_{b}=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$, where $I_{\chi\to b+u}$ is the indicator random variable which takes the value 11 if $\chi\in C_{u}(b)$ and 0 otherwise.
@@
$\mathbb{E}[Y_{b}]=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\Pr[\chi\in C_{u}(b)]=\frac{1}{\Delta^{t}}\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\geq\frac{\epsilon^{2}}{2\Delta^{t}},$ (6.12)
where the second equality follows from the fact that $\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\geq(\frac{\epsilon}{\sqrt{2}})^{2}$ and from Theorem 6.10 (part (1)) where we show that $\Pr[\chi\in C_{u}(b)]=\frac{1}{\Delta^{t}}$.
Before we apply Chebyshev’s inequality to upper bound the probability that a bucket has small weight we upper bound the variance of $Y_{b}$. We can write $Y_{b}=\sum_{\chi\in A_{2}}Y_{\chi,b}$, where $Y_{\chi,b}=|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$. Note that $0\leq Y_{b}\leq\tau$, $\mathbb{E}[Y_{b}]>0$ (by Eq. 6.12) and $\mathsf{cov}(Y_{\chi,b},Y_{\chi^{\prime},b})=0$ (by Theorem 6.10, where $\chi,\chi^{\prime}\in A_{2}$). So, by Lemma 6.11,
@@
$\Pr[Y_{b}\leq(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]]$ $\leq\frac{\operatorname{Var}[Y_{b}]}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]^{2}}\leq\frac{\tau\mathbb{E}[Y_{b}]}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]^{2}}=\frac{\tau}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]}\leq\frac{\tau\cdot 2\Delta^{t}}{\epsilon^{\prime 2}\cdot\epsilon^{2}}.$
Since the total number of buckets is $\Delta^{t}$, therefore the expected number of buckets with weight $\leq(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $\Delta^{t}\cdot\frac{\tau\cdot 2\Delta^{t}}{\epsilon^{\prime 2}\epsilon^{2}}$. Thus by Markov’s inequality, with probability at least $\frac{9}{10}$, the number of buckets with weight more that $(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $10\cdot\frac{\tau\cdot 2\Delta^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot\Delta^{t}$.
@@
$(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]\geq\frac{1}{20}\cdot\frac{\epsilon^{2}}{2\Delta^{t}}=\frac{\epsilon^{2}}{40\Delta^{t}}\geq\tau.$
@@
It thus follows that with probability at least $\frac{9}{10}$, the number of buckets with weight more that $(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $10\cdot\frac{\tau\cdot 2\Delta^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot\Delta^{t}$.
@@
We now finish by showing that $10\cdot\frac{\tau\cdot 2\Delta^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot\Delta^{t}\leq\Delta^{t}-(s+1)$.
$10\cdot\frac{\tau\cdot 2\Delta^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot\Delta^{t}\leq\Delta^{t}-(s+1)$
@@
$\Longleftarrow\tau\leq\frac{\epsilon^{2}}{\Delta^{t}}\cdot\biggl{(}1-\frac{s+1}{\Delta^{t}}\biggr{)}\cdot\left(\frac{19}{20}\right)^{2}\cdot\frac{1}{20}$
$\Longleftarrow\tau\leq\frac{\epsilon^{2}}{\Delta^{t}}\cdot\frac{1}{2}\cdot\frac{1}{20}=\frac{\epsilon^{2}}{40\Delta^{t}}.$